The perimeter of a rectangle is 54 feet and its area is 180 square feet. What are the dimensions of the rectangle?
Answer by edjones(8007)

a) 2(L+w)=54
b) Lw=180
.
a)
2L+2w=54
2L=54-2w
L=27-w
b)
w(27-w)=180
27w-w^2=180
w^2-27w+180=0
(w-12)(w-15)=0
w=12', L=15'
.
Ed